# Reinforcement Fine-Tuning of Large Language Models

### v1

@@ 4.4 Human-AI Collaborative Annotation
This example demonstrates the human-in-the-loop capability in Trinity-RFT for preference modeling. As illustrated in Listing 5 and Figure 5, our framework integrates Label Studio’s annotation interface with asynchronous data pipelines through four coordinated stages: (1) task generation: auto-creating annotation batches from model rollouts; (2) interactive labeling: providing UI for side-by-side response comparison; (3) quality control: enforcing inter-annotator agreement thresholds; and (4) versioned storage: tracking preference lineage in pre-defined fields like those in DPODataModel.
@@
Listing 5: Configuration for human preference annotation.
@@
4.5 Trinity-Studio
Trinity-Studio provides visual interaction for the core capabilities of Trinity-RFT, designed to bridge the gap between system complexity and user accessibility. As shown in Figure LABEL:fig:trinity-studio:dashboard, its three integrated modules, “Training Portal”, pgAdmin”, and “Label Studio”, form a cohesive interface that directly supports the active data-reward co-shaping and human-AI collaboration capabilities established in Section 3.3.
  • “Training Portal” (Figure LABEL:fig:trinity-studio:training) implements configuration-to-execution procedures through declarative YAML editing with live validation and auto completion. This design aligns with Trinity-RFT’s decoupled architecture, enabling users to optionally specify dataset paths, processing and training parameters, meanwhile ensuring schema-aware validation against pre-defined data schema constraints (Section B.1). The live validation mechanism prevents misconfigurations that could disrupt the explorer-trainer coordination described in Section 3.1. Furthermore, the integration of runtime metrics with tools like Wandb/TensorBoard directly helps the active data optimization feature by surfacing signals such as difficulty distribution drifts and diversity metrics mentioned in Section 4.3. This transparency ensures that users can monitor how data curation strategies impact RFT performance in real time.
  • “pgAdmin” (Figure LABEL:fig:trinity-studio:pgadmin) reflects Trinity-RFT’s end-to-end data transformation capabilities by providing a visual panel for PostgreSQL-based storage (Section 3.3.1). This design benefits the versioned data lineage requirements of RFT, particularly for scenarios involving asynchronous training (Section 3.3.2). With intuitive SQL query builders, users can easily adjust schema, audit training experiences and human annotation batches with fine-grained precision. This capability is valuable for rapid validation of active learning policies by cross-referencing training outcomes with metadata (e.g., difficulty scores and staleness in asynchronous mode).
  • “Label Studio” page (Figure LABEL:fig:trinity-studio:label-studio) operationalizes Trinity-RFT’s bi-directional human-AI collaboration capability (Section 3.3.3). Utilizing the provided task polling and atomic batch commit mechanisms, users can annotate the data or experiences directly, allowing an asynchronous way to involve human feedback and to dynamically influence data curation.
By unifying these capabilities in a single UI, Trinity-Studio reduces the cognitive load of managing complex RFT procedures. For example, a researcher tuning a math reasoning task could use the Training Portal to adjust difficulty scoring parameters, view the resulting distribution shifts in the pgAdmin module, and then validate human annotators’ preferences in the Label Studio page. This end-to-end visibility is critical for debugging and iterating on RFT strategies, particularly when balancing tradeoffs between quality, diversity, and reward shaping. Moreover, this visual operation complements the programmatic APIs of Trinity-RFT while maintaining full compatibility with CLI procedures. We implement Trinity-Studio with Singe-Spa framework [27]. The modular architecture enables custom view development through JavaScript plugins and flexible extensions for general-purpose usage.
@@ 4.5 Trinity-Studio
5 Conclusion
This report has provided an overview of Trinity-RFT, a general-purpose, flexible, scalable and user-friendly framework for reinforcement fine-tuning of large language models. Trinity-RFT offers a path into “the era of experience” [26], by supporting applications in diverse scenarios with complex agent-environment interaction, and serving as a unified platform for exploring advanced RL paradigms.
@@ Acknowledgements
Trinity-RFT is built upon many excellent open-source projects, including but not limited to: verl [24] and PyTorch’s FSDP [5] for LLM training; vLLM [10] for LLM inference; Data-Juicer [2] for data processing functionalities; AgentScope [6] for agentic workflow; and Ray [14] for distributed systems. We have also drawn inspiration from RL/RFT projects like OpenRLHF [8], TRL [31], ChatLearn [1], and TinyZero [19].
@@ Appendix A Off-Policy Reinforcement Learning Algorithms
In the rest of this section, we present some of our findings about off-policy RL (with a focus on the bandit setting) through the lens of online policy mirror descent (OPMD). At the end of this section, we arrive at the surprising conclusion that the standard policy gradient, weighted by a coefficient and using the group mean reward as the baseline, can be a feasible direction for updating the policy even in off-policy settings (while the standard theory of policy gradient only holds for on-policy settings). This has been validated empirically in our exploratory experiments during the development of Trinity-RFT.
@@ A.3 OPMD: an Embarrassingly Simple Variant
As a heuristic, we simply add a regularization term (denoted by $g$) to the above objective when additional regularization with respect to a fixed policy, e.g., a SFT model $\pi_{\textsf{sft}}$, is desired:
@@
Appendix B Additional Details for Data Pipelines
B.1 Illustration of Typical Dataset Classes and Schemas
The dataset shaping process described in Section 3.3 is supported by PostgreSQL persistence with comprehensive schema designs; one example can be found in Listing 6. Key model classes include RftDatasetModel (core dataset storage with lineage tracking), ExperienceModel (which stores serialized experiences with priority scoring), DPODataModel and SFTDataModel (specialized storage for DPO and SFT, respectively).
1-- RftDataset table structure
2CREATE TABLE rft_dataset (
3 id SERIAL PRIMARY KEY, -- Unique record identifier
4 consumed_cnt INTEGER DEFAULT 0, -- Usage counter for sampling
5 prompt TEXT, -- Input instruction
6 response TEXT, -- Model-generated response
7 reward FLOAT, -- Training signal value
8 quality_score FLOAT DEFAULT 0.0, -- Quality assessment metric
9 difficulty_score FLOAT DEFAULT 0.0, -- Complexity estimation
10 diversity_score FLOAT DEFAULT 0.0, -- Response uniqueness measure
11 priority FLOAT DEFAULT 0.0 -- Sampling preference weight
12);
Listing 6: Schema definition for RftDataset PostgreSQL table.
B.2 Illustration of Data Processing Service
The architecture presented in Section 3.3 integrates Data-Juicer capabilities through service-oriented design. Typically, the DJ-service lifecycle covers three stages: initialization via scripts/start_servers.py; configuration through dj_config_path or natural language (dj_process_desc); and execution handling dataset loading, cleaning, priority scoring, and database export. Moreover, we provide DJExecutor to bridge Data-Juicer capabilities through config translation and execution sandboxing, and dynamic termination via _check_clean_targets() with melt protection.
B.3 An Example of Active Data Optimization Algorithms
The DataActiveIterator serves as the core engine for data-model co-evolution with multi-dimensional utility scoring. It provides extensible interfaces for flexible training prior injection, e.g., as demonstrated in Algorithm 1.
Algorithm 1 Multi-Dimensional Sample Selection
0:  Dataset $D$, weights $w_{q},w_{d},w_{k},w_{f}$
0:  Selected subset $D^{\prime\prime}$
1:  $(q,d,k)\leftarrow\text{DataCleaner}.\text{Process}(D)$ Compute quality ($q$), diversity ($d$), difficulty ($k$)
2:  $U_{i}\leftarrow w_{q}q_{i}+w_{d}d_{i}+w_{k}k_{i}-w_{f}f_{i}$ Combine $q$, $d$, $k$, frequency ($f$)
3:  $D^{\prime}\leftarrow\{s_{i}\in D\mid U_{i}>\theta_{\text{min}}\}$ Threshold filtering
4:  $D^{\prime\prime}\leftarrow\text{Top-K}(D^{\prime},K)$ Priority selection
5:  $\text{ExperienceCache}.\text{Update}(D^{\prime\prime})$ Update version-controlled cache
6:  $\text{LineageTracker}.\text{Record}(D^{\prime\prime})$ Track data provenance
7:  return $D^{\prime\prime}$

### v2

@@ 3.5 RFT with Human in the Loop
This example demonstrates the human-in-the-loop capability in Trinity-RFT for preference modeling. As illustrated in Listing 6 and Figure 15, our framework integrates Label Studio’s annotation interface with asynchronous data pipelines through four coordinated stages: (1) task generation: auto-creating annotation batches from model rollouts; (2) interactive labeling: providing UI for side-by-side response comparison; (3) quality control: enforcing inter-annotator agreement thresholds; and (4) versioned storage: tracking preference lineage in pre-defined fields like those in DPODataModel.
@@
Listing 6: Configuration for human preference annotation.
@@
3.6 Low-Code Usage and Development with Trinity-Studio
(a) Trinity-Studio dashboard.
(b) Start training on the “Training Portal” page.
(c) Manage data on the “pgAdmin” page.
(d) Process data on the “Label Studio” page.
@@ 3.6 Low-Code Usage and Development with Trinity-Studio
Trinity-Studio provides visual interaction for the core capabilities of Trinity-RFT, designed to bridge the gap between system complexity and user accessibility. As shown in Figure 16(a), its three integrated modules — “Training Portal”, “pgAdmin”, and “Label Studio” — form a cohesive interface that supports low-code usage and development with Trinity-RFT, and make it easy to monitor and track the full RFT pipeline with transparency.
  • “Training Portal” (Figure 16(b)) implements configuration-to-execution procedures through declarative YAML editing with auto completion and live validation that prevents misconfigurations. Furthermore, the integration of runtime metrics with tools like Wandb/TensorBoard directly helps the active data optimization feature by surfacing signals such as difficulty distribution drifts and diversity metrics mentioned in Section 3.4. This transparency ensures that users can monitor how data curation strategies impact RFT performance in real time.
  • “pgAdmin” (Figure 16(c)) reflects Trinity-RFT’s end-to-end data transformation capabilities by providing a visual panel for PostgreSQL-based storage. This design benefits the versioned data lineage requirements of RFT, particularly for scenarios involving asynchronous training (Section 2.3.3). With intuitive SQL query builders, users can easily adjust schema, audit training experiences and human annotation batches with fine-grained precision. This capability is valuable for rapid validation of active learning policies by cross-referencing training outcomes with metadata (e.g., difficulty scores and staleness in asynchronous mode).
  • “Label Studio” page (Figure 16(d)) operationalizes Trinity-RFT’s bi-directional human-AI collaboration capability (Section 2.3.4). Utilizing the provided task polling and atomic batch commit mechanisms, users can annotate the data or experiences directly, allowing an asynchronous way to involve human feedback and to dynamically influence data curation.
By unifying these capabilities in a single UI, Trinity-Studio reduces the cognitive load of managing complex RFT procedures. For example, a researcher tuning a math reasoning task could use the Training Portal to adjust difficulty scoring parameters, view the resulting distribution shifts in the pgAdmin module, and then validate human annotators’ preferences in the Label Studio page. This end-to-end visibility can be useful for debugging and iterating RFT strategies, and complements the programmatic APIs of Trinity-RFT while maintaining full compatibility with CLI procedures.
We implement Trinity-Studio with the Singe-Spa framework [32]. The modular architecture enables custom view development through JavaScript plugins and flexible extensions for general-purpose usage.
4 Conclusion and Next Steps
We have presented Trinity-RFT, a general-purpose, flexible, scalable and user-friendly framework for reinforcement fine-tuning of large language models. Trinity-RFT offers a path into “the era of experience” [31], by supporting applications in diverse scenarios with complex agent-environment interaction, and serving as a unified platform for exploring advanced methodologies in each stage of the complete RFT pipeline, at both macroscopic and microscopic levels.
Further development of Trinity-RFT includes enhancing support for generic multi-step agentic workflows, incorporating more advanced RL algorithms (especially off-policy or asynchronous ones), making the choices of hyperparameters more adaptive and less reliant on manual tuning, augmenting data pipelines with smarter sampling strategies and data processing operations, and conducting more thorough experiments and evaluations with Trinity-RFT.
@@ Acknowledgements
Trinity-RFT is built upon many excellent open-source projects, including but not limited to: verl [29] and PyTorch’s FSDP [7] for LLM training; vLLM [14] for LLM inference; Data-Juicer [2] for data-related functionalities; AgentScope [10] for agentic workflow; and Ray [19] for distributed runtime.
@@ Appendix A Off-Policy Reinforcement Learning Algorithms
In the rest of this section, we present some of our findings about off-policy RL (with a focus on the bandit setting) through the lens of online policy mirror descent (OPMD). At the end of this section, we arrive at the surprising conclusion that the standard policy gradient, weighted by a coefficient and using the group mean reward as the baseline, can be a feasible direction for updating the policy even in off-policy settings (while the standard theory of policy gradient only holds for on-policy settings). This has been validated empirically in our exploratory experiments during the development of Trinity-RFT  [^14] https://github.com/modelscope/Trinity-RFT/blob/main/docs/sphinx_doc/source/tutorial/example_reasoning_advanced.md.
@@ A.3 OPMD: an Embarrassingly Simple Variant
Remark 2.
As a heuristic, when additional regularization with respect to a fixed policy (e.g., a SFT model $\pi_{\textsf{sft}}$) is desired, we simply add a regularization term (denoted by $g$) to the above objective: